Grade 12Organic Chemistry

define current sensitivity of galvanometer. write any two factors on which it depends.

sheikh
7 Years agoGrade 12

1 Answer

jyoti bhatia
7 Years ago
A galvanometer is said to be sensitive, if it gives a large deflection, even when a small current is passed through it or when a small voltage is applied across its coil. 
     There are two types of  sensitive galvanometer depends as follows :- 
1. current sensitivity :-  current sensitivity is defined as the deflection  produced in the galvanometer on passing a unit current through its coil,
   e.g.,Is =  θ/I = NBA/K
where, K is a torsion constant for galvanometer.
            N is the number of turns
            B is the magnetic field
            θ is the angle by which torque displace the coil.
2. voltage sensitivity :-  voltage sensitivity is defined as the deflection produced in a galvanometer when a unit potential difference is applied across its end.
   e.g., Vs =  θ/V = θ/IR = NBA/KR 
Now you can say that sensitivity of galvanometer depends upon :-
  1. it is directly proportional to the area of coil.
  2. it is directly proportional to the magnetic field.
  3. it is inversely proportional to the torsion constant.
  4. it is directly proportional to number of turns.